# Beyond Perplexity: Multi-dimensional Safety Evaluation of LLM Compression

### v2

@@ Abstract
Large language models (LLMs) are increasingly deployed in real-world scenarios with the help of recent model compression techniques. Such momentum towards local deployment means the use of compressed LLMs will widely impact a large population. However, prior analysis works often prioritize on preserving perplexity which is a direct analogy to training loss. The impact of compression method on other critical aspects of model behavior, particularly safety, still calls for a systematic assessment. To this end, we investigate the impact of model compression on four dimensions: (1) degeneration harm, i.e., bias and toxicity in generation; (2) representational harm, i.e., biases in discriminative tasks; (3) dialect bias; (4) language modeling and downstream task performance. We cover a wide spectrum of LLM compression techniques, including unstructured pruning, semi-structured pruning and quantization. Our analysis reveals that compression can lead to unexpected consequences. Although compression may unintentionally remedy LLMs’ degeneration harm, it can still exacerbate on the representational harm axis. Moreover, there is a divergent impact on different protected groups as the compression rate grows. Finally, different compression methods have drastically different safety impacts, e.g., quantization mostly preserves bias while pruning degrades quickly. Our findings underscore the importance of integrating safety assessments into the development of compressed LLMs to ensure their reliability across real-world applications. [^1] Our full results are available here: https://github.com/zhichaoxu-shufe/Beyond-Perplexity-Compression-Safety-Eval
@@ 1 Introduction
Large language models (e.g., Gemini et al., 2023; Achiam et al., 2023) are remarkably performant across various tasks; they have been deployed not only as intelligent assistants such as ChatGPT, but also in high-stake scenarios such as psychology Demszky et al. (2023) and medical diagnosis Saab et al. (2024). The sensitivity of such applications necessitates evaluating them across multiple dimensions, including accuracy, robustness, and other factors (Gupta et al., 2023; Liang et al., 2023).
Despite LLM’s potential usefulness, their high computational costs render local deployments difficult (cf. Zhu et al., 2023; Chien et al., 2023). Consequently, there has been a surge of interest in compression methods that convert LLMs into compact models for efficient storage and inference by reducing their latency as well as memory footprint (e.g., Sun et al., 2024; Frantar and Alistarh, 2023; Lin et al., 2024; Ma et al., 2023; Frantar et al., 2022). Pruning algorithms like SparseGPT Frantar and Alistarh (2023) and Wanda Sun et al. (2024) can substantially reduce the number of active LLM parameters without compromising perplexity. Similarly, quantization methods (e.g., Lin et al., 2024; Dettmers et al., 2022; Frantar et al., 2022) can reduce the memory footprint of LLMs by reducing bit-precision during inference without significantly impacting perplexity.
Model compression methods largely focus on ensuring that the perplexity of the compressed models does not deteriorate. However, solely relying on perplexity as a performance metric is insufficient. For example, compressing large language models by a small fraction (e.g., a 20% reduction) may result in minimal changes in perplexity, but can lead to significant degradation in performance on downstream tasks (Hong et al., 2024; Yin et al., 2023). More importantly, there is a lack of a systematic evaluation of how compression affects an LLM along safety dimensions, such bias, toxicity and truthfulness. This gap prevents the successful application of such compressed LLMs in sensitive applications.
In this work, we argue that usage costs and data sharing restrictions will mean that local deployments of compressed LLMs are more likely to impact a larger population. Given the potential widespread use of compressed LLMs, we ask: Are compressed LLMs not only accurate, but also safe? To this end, we conduct a multi-faceted evaluation of compressed LLMs, including: (1) evaluating its degeneration harm, i.e. model’s biased and toxic behaviors in the generated text; (2) evaluating its representational harm, which arises when language models are deployed for discriminative tasks; (3) evaluating how LLM compression affects dialect bias, and (4) the impact of compression on model’s language modeling capabilities and downstream task performances. We cover a wide spectrum of compression methods, including unstructured pruning, semi-structured pruning and quantization. Some of our key findings are:
@@
  • Quantization methods mostly preserve model’s bias, toxicity and performance at a moderate compression rate (e.g. 50%), while pruning methods experience significant degradation at the same compression rate.
@@ 2.1 Potential Harms by LLMs
Representational Harm  Different form degeneration harm, which manifests during text generation, representational harm arises when LLMs are deployed for discriminative tasks, such as text classification or representation learning (Wang et al., 2022; Crawford, 2017). Existing works on measuring representational harm primarily examine model’s behaviors with respect to various protected characteristics like religion, gender, etc. through under-specified questions Parrish et al. (2022); Li et al. (2020). For instance, when asked about which pronouns are more likely to be associated with computer programmers, BERT-style question answering models prefer male pronouns to female pronouns, despite the gender of the occupation not being specified in the question’s context (Li et al., 2020). We provide experimental details for measuring these two types of harms in Sec. 4.
@@
Our goal is to evaluate the safety of compressed LLMs. Notable compression techniques include network pruning LeCun et al. (1989); Hassibi et al. (1993), distillation Sanh et al. (2019), quantization Dettmers et al. (2022); Frantar et al. (2022); Lin et al. (2024) and low-rank approximation Xu et al. (2023); Lan et al. (2019).
@@ 3.1 Compression Algorithms and Ratios
We also include representative post-training quantization methods—LLM.int8() Dettmers et al. (2022) [^4] Commonly referred to as BitsAndBytes 8-bit quantization, GPTQ Frantar et al. (2022) and Activation-aware Weight Quantization (AWQLin et al. (2024). Inputs and weights in LLM.int8() are multiplied in 8-bit and quantized to Int8 before being dequantized back to 16-bits. GPTQ is a layer-wise quantization technique based on approximated second-order information towards minimum accuracy loss on the calibration set. AWQ reserves some salient weights in 16-bits while quantizing other weights to 4-bits without significant performance degradation. Table 1 compares the compression methods, and we show additional technical details in Appx. B .
@@ 4 Degeneration Harm & Representational Harms
(a) Evaluation results of Llama-2-13B on language modeling ($\downarrow$), toxicity ($\downarrow$) and bias datasets ($\downarrow$). We notice model-based evaluation metrics are sensitive to generation quality, e.g. % negative regard decreases as perplexity increases.
@@
Quantized 13B models are on par or better than uncompressed 7B models. The 50% quantized Tülu-2-13B model with LLM.int8() achieves 56.7% and 55.6% on MMLU and TruthfulQA datasets, compared to the original Tülu-2-7B model’s 55.8% and 32.3%. Note that these two models are roughly equal in terms of the raw memory they need. In terms of language modeling, 50% quantized Llama-2-13B model achieves 4.92 perplexity on WikiText-2 compared to Llama-2-7B’s 5.47. On the other hand, 50% pruned Tülu-2-13B with GBLM pruning only achieves 51.3% and 44.4% on MMLU and TruthfulQA, respectively. This suggests that under same compression rate, quantization performs better than pruning.
@@ 6 How Does Compression Affect Different Dialects of English?
Different prior works have studied dialect biases for language models (Blodgett et al., 2016, 2020; Joshi et al., 2024; Lent et al., 2021, inter alia). Notably, Hofmann et al. (2024) highlight that LLMs may embody covert racism in the form of dialect prejudice. In this section, we study how compression affects language models’ dialect biases. Specifically, we focus on African American English (AAE) versus "standard" English. We use two paired datasets for this evaluation: (1) Twitter AAE dataset Blodgett et al. (2020), consisting of balanced sets of tweets classified as African American or White-aligned English; and (2) AAE literature dataset [^5] https://github.com/jazmiahenry/aave_corpora versus Dolma books subset Soldaini et al. (2024). The first comparison focuses on social media posts while the second comparison focuses on public domain of books. We show the detailed statistics of these datasets in Table 5. We evaluate the change of perplexity of pruned language models on these corpora. This comparison provides us insights into how different pruning methods and pruning ratios affect the language model’s dialect biases.
We show the results with Llama-2-13b base model in Fig. 3. The full results are in Sec. D.3. We make three key observations: (1) The pre-trained language model has a dialect bias. It has a lower perplexity on standard English book text or social media posts by the White ethic group, compared to their AAE counterparts. (2) Model compression maintains the language model’s dialect biases. The perplexity of both AAE and "standard" English increases as the compression ratio increases, but the margin does not reduce. This is true for both pruning and quantization methods. (3) The perplexity of even a heavily compressed model (at 50% pruning ratio) on "standard" English is better than that of the uncompressed model on African American English.
The impact of the last observation can be illustrated by mapping model size to monetary cost of inference; larger models cost more. The largest (i.e. uncompressed) model is double the size of the 50% compressed model, but the former has worse perplexity on AAE than the latter on standard English. As language models are increasingly becoming our interfaces to data and compute, this means that a speaker of White-aligned English can receive "better service" in their native dialect, but pay only half the price as an AAE speaker seeking to interact in their native dialect.
7 The Impact of Supervised Fine-tuning
@@ 8 Conclusions and Recommendations
In this work, we presented a comprehensive evaluation on the safety of LLM compression techniques. We took a systematic perspective on the safety dimension by investigating multiple aspects, including degeneration harm, representational harm, and dialect biases. Our safety evaluation, along with downstream task performances, revealed that model compression can lead to a series of unexpected results, including the divergent changes of different harms and downstream task scores. Our findings highlight the need for a nuanced understanding of how compression affects LLM behavior. To future compression works, we hereby recommend the following: 1) do not just measure perpexity or safety, always use both; 2) aggregated metrics for safety can hide the nuanced movement across different protected groups and dialects, and it is imperative to conduct fine-grained evaluation for compressed LLMs with regard to each individual protected groups and dialects.
9 Limitations
@@
The compression algorithms and representational harm evaluations require access to model’s parameters and logits, which are not available for certain proprietary models such as GPT-4 Achiam et al. (2023) and Gemini Gemini et al. (2023).
@@ Acknowledgements
We would thank members of UtahNLP for their constructive feedback. This material is based upon work supported in part by NSF under grants 2007398, 2217154, 2318550, 2205418, and 2134223. Ashim Gupta is supported by the Bloomberg Data Science Ph.D. Fellowship. Oliver Bentham is supported by the NSF CISE Graduate Fellowships under Grant No. G-2A-063. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsers.
@@ A.1.1 Generative Datasets
  • BOLD Dhamala et al. (2021) [^12] https://huggingface.co/datasets/AlexaAI/bold is a bias dataset that contains 7,201 prompts covering 5 different dimensions: profession, gender, race, religious ideology and political biology. Each dimension includes several groups. We follow (Touvron et al., 2023) to study how the sentiment in model generations may vary with groups. We evaluate the sentiment w.r.t. each group with VADER classifier Hutto and Gilbert (2014), a ruled-based sentiment classifier adopted in the Llama-2 Touvron et al. (2023)’s evaluation.
@@ A.1.2 Representational Bias Datasets
    Here, the score $\eta(x,a)$ is defined in Li et al. (2020) with $x$ denotes a subject and $a$ an attribute. We refer the reader to Li et al. (2020) for details of derivation of the metric $\eta(x,a)$.

### v3

@@ Abstract
Increasingly, model compression techniques enable large language models (LLMs) to be deployed in real-world applications. As a result of this momentum towards local deployment, compressed LLMs will interact with a large population. Prior work on compression typically prioritize preserving perplexity, which is directly analogous to training loss. The impact of compression method on other critical aspects of model behavior — particularly safety — requires systematic assessment. To this end, we investigate the impact of model compression along four dimensions: (1) degeneration harm, i.e., bias and toxicity in generation; (2) representational harm, i.e., biases in discriminative tasks; (3) dialect bias; and (4) language modeling and downstream task performance. We examine a wide spectrum of LLM compression techniques, including unstructured pruning, semi-structured pruning, and quantization. Our analysis reveals that compression can lead to unexpected consequences. Although compression may unintentionally alleviate LLMs’ degeneration harm, it can still exacerbate representational harm. Furthermore, increasing compression produces a divergent impact on different protected groups. Finally, different compression methods have drastically different safety impacts: for example, quantization mostly preserves bias while pruning degrades quickly. Our findings underscore the importance of integrating safety assessments into the development of compressed LLMs to ensure their reliability across real-world applications. [^1] Our implementation and results are available here: https://github.com/zhichaoxu-shufe/Beyond-Perplexity-Compression-Safety-Eval
@@ 1 Introduction
Large language models (e.g., Gemini et al., 2023; Achiam et al., 2023) are remarkably performant across various tasks; they have been deployed not only as intelligent assistants, but also in high-stake scenarios such as psychology Demszky et al. (2023) and medical diagnosis Saab et al. (2024). The sensitivity of such applications necessitates evaluating them across multiple dimensions, including accuracy, robustness, and other factors (Gupta et al., 2023b; Liang et al., 2023).
Despite potential usefulness, high computational costs render local LLM deployments difficult (cf. Zhu et al., 2023; Chien et al., 2023). Consequently, there has been a surge of interest in compression methods that convert LLMs into compact models for efficient storage and inference by reducing their latency as well as memory footprint (e.g., Sun et al., 2024; Frantar and Alistarh, 2023; Lin et al., 2024; Ma et al., 2023; Frantar et al., 2022). Pruning algorithms like SparseGPT Frantar and Alistarh (2023) and Wanda Sun et al. (2024) can substantially reduce the number of active LLM parameters without compromising perplexity. Similarly, quantization methods (e.g., Lin et al., 2024; Dettmers et al., 2022; Frantar et al., 2022) can reduce the memory footprint of LLMs by reducing bit-precision during inference without significantly impacting perplexity.
Model compression methods primarily focus on ensuring that the perplexity of the compressed models does not deteriorate. However, solely relying on perplexity as a performance metric is insufficient. For example, compressing large language models by a small fraction (e.g., a 20% reduction) may result in minimal changes in perplexity, but can lead to significant degradation in performance on downstream tasks (Hong et al., 2024; Yin et al., 2023). More importantly, there is a lack of systematic evaluation of how compression affects an LLM along safety dimensions, such as bias, toxicity, and truthfulness.
In this work, we argue that usage costs and data sharing restrictions will mean that local deployments of compressed LLMs are more likely to impact a larger population. Given their potential widespread use, we ask: Are compressed LLMs not only accurate, but also safe? To this end, we conduct a multi-faceted evaluation of compressed LLMs, including: (1) evaluating its degeneration harm, i.e. toxicity and bias in model generated text; (2) evaluating its representational harm, which arises when language models are deployed for discriminative tasks; (3) evaluating how LLM compression affects dialect bias, and (4) the impact of compression on model’s language modeling capabilities and downstream task performance. We cover a wide spectrum of compression methods, including unstructured pruning, semi-structured pruning and quantization. Some of our key findings are:
@@
  • Pretrained language models have dialect biases, and model compression maintains such biases.
  • Quantization methods mostly preserve model’s bias, toxicity and performance at a moderate compression rate (e.g. 50%), while pruning methods show significant degradation at the same compression rate.
@@ 2.1 Potential Harms by LLMs
Representational Harm  Different form degeneration harm, which manifests during text generation, representational harm arises when LLMs are deployed for discriminative tasks, such as text classification (Wang et al., 2022; Crawford, 2017). [^2] Barocas et al. (2023) mention stereotype perpetuation and cultural denigration as examples of representational harms, and argue that they “occur when systems reinforce the subordination of some groups along the lines of identity — race, class, gender, etc. [They] have long-term effects, and resist formal characterization.” In our experiments, we use the BBQ and UnQover evaluations to focus on the stereotype perpetuation aspect, and evaluate the extent to which language model reinforces stereotypes against protected groups. Existing works on measuring representational harm primarily examine models’ behaviors with respect to various protected characteristics such as religion and gender via under-specified questions Parrish et al. (2022); Li et al. (2020). For instance, when asked which pronouns are more likely to be associated with computer programmers, BERT-style question answering models prefer male pronouns to female pronouns, despite the gender of the occupation not being specified in the question’s context (Li et al., 2020). We provide experimental details for measuring these two types of harms in Sec. 4.
@@
Our goal is to evaluate the safety of compressed LLMs. Notable compression techniques include network pruning LeCun et al. (1989); Hassibi et al. (1993); Xia et al. (2022, 2024), distillation Sanh et al. (2019), quantization Dettmers et al. (2022); Frantar et al. (2022); Lin et al. (2024); Zhang and Shrivastava (2024) and low-rank approximation (Gupta et al., 2024; Xu et al., 2023; Lan et al., 2019, inter alia).
@@
2.3 Prior Works on LLM Compression Evaluation
A few recent works have attempted to tackle the problem of safety evaluation of LLM compression. For example, Ramesh et al. (2023) evaluate how different compression methods affect language model’s fairness dimensions, but the experiments are restricted to moderate-sized, encoder-only models. Jaiswal et al. (2023) highlight the problem of using perplexity as the standalone evaluation metric and underscore the importance of more comprehensive evaluations, yet their experiments are restricted to performance dimensions of compressed LLMs. Different from Hong et al. (2024) which evaluates "trustworthiness" of compressed LLMs as an aggregated score, in this work we attempt to conduct a fine-grained, multifaceted safety evaluation of compressed LLMs, with particular attention to disparities in how model compression affects different protected groups.
@@ 3.1 Compression Algorithms and Ratios
We also include representative post-training quantization methods — LLM.int8() Dettmers et al. (2022), GPTQ Frantar et al. (2022) and Activation-aware Weight Quantization (AWQLin et al. (2024). Inputs and weights in LLM.int8() are multiplied in 8-bit and quantized to Int8 before being dequantized back to 16-bits. GPTQ is a layer-wise quantization technique based on approximated second-order information towards minimum accuracy loss on the calibration set. AWQ reserves some salient weights in 16-bits while quantizing other weights to 4-bits without significant performance degradation. Table 1 compares the compression methods, and we show additional technical details in Appx. B.
@@ 3.2 Safety Evaluation Dimensions
(a) Evaluation results of Llama-2-13B on language modeling ($\downarrow$), toxicity ($\downarrow$) and bias datasets ($\downarrow$). We notice model-based evaluation metrics are sensitive to generation quality, e.g. % negative regard decreases as perplexity increases. Note that from 30% compression rate, all four pruning methods have statistically significantly higher perplexity compared to uncompressed models (paired student T-Test at 0.05 significance level).
@@ 4 Degeneration Harm & Representational Harms
Quantized 13B models are on par or better than uncompressed 7B models. The 50% quantized Tülu-2-13B model with LLM.int8() achieves 56.7% and 55.6% on MMLU and TruthfulQA datasets, compared to the original Tülu-2-7B model’s 55.8% and 32.3%. Note that these two models are roughly equal in terms of the GPU memory they require for inference. In terms of language modeling, 50% quantized Llama-2-13B model achieves 4.92 perplexity on WikiText-2 compared to Llama-2-7B’s 5.47. On the other hand, 50% pruned Tülu-2-13B with GBLM pruning only achieves 51.3% and 44.4% on MMLU and TruthfulQA, respectively. This suggests that under same compression rate, quantization performs better than pruning.
@@ 6 How Does Compression Affect Different Dialects of English?
Different prior works have studied dialect biases for language models (Blodgett et al., 2016, 2020; Joshi et al., 2024; Lent et al., 2021, inter alia). Notably, Hofmann et al. (2024) highlight that LLMs may encode systemic racial biases via dialect prejudice. In this section, we study how compression affects language models’ dialect biases. Specifically, we focus on African American English (AAE) versus "standard" English. We use two paired datasets for this evaluation: (1) the Twitter AAE dataset Blodgett et al. (2020), consisting of balanced sets of tweets classified as African American or White-aligned English; (2) the AAE literature dataset [^5] https://github.com/jazmiahenry/aave_corpora versus Dolma books subset Soldaini et al. (2024). The first comparison focuses on social media posts while the second comparison focuses on public domain books, representing (typically) copy-edited text. We provide the detailed statistics of these datasets in Table 5. We evaluate the change of perplexity of compressed language models on these corpora. This comparison provides us insights into how different compression methods and compression ratios affect the language model’s dialect biases.
We show the results with Llama-2-13B base model in Fig. 3. The full results are in Sec. D.3. We make three key observations: (1) The pre-trained language model has a dialect bias. It has a lower perplexity on standard English book text or social media posts, compared to their African American English counterparts. (2) Model compression maintains the language model’s dialect biases. The perplexity of both AAE and “standard” English increases as the compression ratio increases, but the margin does not reduce. This is true for both pruning and quantization methods. (3) Even a heavily compressed model (at 50% pruning ratio) has better perplexity on "standard" English than the uncompressed model on African American English. Notwithstanding the difficulty of the selection of the perplexity evaluation dataset and the underlying phenomenon of dialect bias, our conclusions remain valid because of the significantly worse perplexity of AAE dialect with the uncompressed model.
The impact of the last observation can be illustrated by mapping model size to monetary cost of inference; larger models cost more. The largest (i.e. uncompressed) model is double the size of the 50% compressed model, but the former has worse perplexity on AAE than the latter on standard English. As language models are increasingly becoming our interfaces to data and compute, this means that a speaker of White-aligned English can receive “better service” in their native dialect, but pay only half the price as an AAE speaker seeking to interact in their native dialect.
@@ 8 Conclusions and Recommendations
In this work, we presented a comprehensive evaluation on the safety of LLM compression techniques. We systematically investigated multiple aspects of safety, including degeneration harm, representational harm as well as dialect biases. Our safety evaluation, along with downstream task performances, reveals that model compression can lead to a series of unexpected results. Compression may unintentionally remedy an LLM’s degeneration harm, but it can still exacerbate representational harm. In addition, as the compression rate increases, different protected groups are not affected equally. Our findings highlight the need for a nuanced understanding of how compression affects LLM behavior. We conclude with the following recommendations for future LLM compression research: (1) Do not solely evaluate one aspect, perplexity or safety, in isolation. Instead, always measure and report both. (2) Aggregated metrics for safety can hide the nuanced movement across different protected groups and dialects. It is imperative to conduct fine-grained evaluations of compressed LLMs with regard to each individual protected group and dialect.
Limitations
@@
Ethical Considerations
This work studies how model compression affects language model’s safety dimensions, including degeneration harm, representational harm as well as dialect biases. All artifacts used in this work are available for public access with licenses for academic purposes. Given the fact that we conclude compressed models have the same or higher harms, we do not plan to release the compressed models, but we will release detailed implementations and instructions to reproduce the experimental results.
@@ Acknowledgements
We would thank members of UtahNLP for their constructive feedback. This material is based upon work supported in part by NSF under grants 2007398, 2217154, 2318550, 2205418, and 2134223. This research is supported by the National Artificial Intelligence Research Resource (NAIRR) Pilot and the Delta advanced computing and data resource which is supported by the National Science Foundation (award NSF-OAC 2005572). Ashim Gupta is supported by the Bloomberg Data Science Ph.D. Fellowship. Oliver Bentham is supported by the NSF CISE Graduate Fellowships under Grant No. G-2A-063. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsers.
@@ A.1.1 Generative Datasets
  • BOLD Dhamala et al. (2021) [^12] https://huggingface.co/datasets/AlexaAI/bold is a bias dataset that contains 7,201 prompts covering 5 different dimensions: profession, gender, race, religious ideology and political biology. Each dimension includes several groups. We follow Touvron et al. (2023) to study how the sentiment in model generations may vary with groups. We evaluate the sentiment w.r.t. each group with VADER classifier Hutto and Gilbert (2014), a ruled-based sentiment classifier adopted in the Llama-2 Touvron et al. (2023)’s evaluation.
@@ A.1.2 Representational Bias Datasets
    Here, the score $\eta(x,a)$ is defined in Li et al. (2020) with $x$ denotes a subject and $a$ an attribute. We refer the reader to Li et al. (2020) for details of the derivation of the metric $\eta(x,a)$.